Please help solve
Translate to an algebra statement:
Five times the difference of a number and one yields the same result as twice the same number increased by three.

Answer by KnightOwlTutor(293)

The algebraic equation is
5(X-1)=2X+3
Use the distributive property
5X-5=2X+3
Subtract 2X from both sides of the equation
3X-5=3
Add 5 to both sides of the equation
3X=8
Divdie both sides by 8
X=8/3
Replace all X in the above equation to confirm
5(8/3-1)=2(8/3) + 3
replace 1 with 3/3 amd 3 with 9/3
5(8/3-3/3) =2(8/3) + 9/3
5(5/3)=16/3 +9/3
25/3=25/3